# Tidal deformability in neutron stars from a microscopic point of view

### v1

@@ IV.1 General aspects
The tidal deformability is a parameter that quantifies how easily the star is deformed by an external tidal field. Therefore, a largetidal deformability signals a larger and less compact star which deformes easily.
@@
In the Newtonian limit, the perturbing tidal field, $\epsilon_{ij}$, is the second spacial derivative of the external field, resulting in units of inverse length squared [59].The quadrupole moment has units of length cubed. We expect $\lambda_{t}$ to be proportional to the fifth power of the neutron star radius, $R$. Adopting traditional convention, we express the tidal deformability as
@@
where C is the previously defined neutron star’s compactness. It is $\Lambda$ that actually governs the effects of tides on inspiral waveforms VanOe+2017.
Given a proposed equation of state, the tidal deformability of a neutron star of a certain mass can be obtained by computing the metric in the asymptotic regime using the Einstein equations, extracting the corresponding $r$-order terms, and taking their ratio. The ensuing calculation is described in detail in Ref. Hind2008. The result is an expression for the gravitational Love number, $k_{2}=k_{2}(C,y_{R})$, with $y_{R}$ the solution at the star’s surface of the first order differential equation
@@
and then used in Eq.( 10) to determine the tidal deformability.
Figure 3: Left: The dimensionless tidal deformability as a function of the neutron star mass (left) and as a function of the compactness (right). The pink dots and their uncertainties are taken from the Bayesian analysis of Ref. LH19. See text for more details.
@@ IV.2.1 Tidal deformability and tidal Love number
Closely related to $\Lambda$ is the tidal Love number $k_{2}$, related to $\Lambda$ as in Eq. (10). This is shown in Fig. 4 as a function of the neutron star mass and compactness. Often, comparison between relativistic and Newtonian calculations or tabulations of deformation properties for polytropic EoS with different polytropic indices are given in terms of $k_{2}$ rather than $\Lambda$ Hind2008. Our predicted values of $k_{2}$ are shown in Table 2 for a representative case.
@@
In the context of compactness and tidal deformability, it is insightful to look at the pressure profile of neutron stars with different masses, see Fig. 5. As expected, the central pressure is largest (smallest) for the EoS that results in the smallest (largest) radius for the same mass, compare with Fig. 1.
Figure 5: Pressure profile in neutron stars with different masses as given in the inset.
Recent multimessanger constraints can be found in Ref. Huang25, obtained from various methodologies based on variants of the GW170817 data sets. Their most robust constraint is based on a combination of GW-based findings and EoS-independent X-ray constraints, and is reported as
@@
for the dimensionless tidal deformability and radius of the canonical mass neutron star. We underline that this EoS-independent constraint is yet another confirmation that the large radius extracted from the PREX neutron skin measurement Ree21 is unrealistic.
Accounting for both chiral truncation error and the spreading due to different values of $\gamma$, we estimate our ranges for $\Lambda_{1.4}$ and $R_{1.4}$ at N2LO to be
@@
quite consistent with Eq. (15). We underline that the overall uncertainties in Eq. (16) are evaluated at N2LO, which, at this time, is a fully consistent order, see comments in Sec. II.
Other inferred upper limits on neutron star radii and tidal deformabilities from GW170817 rule out stiff equations of state that result in large radii and large tidal deformabilities. Recent estimates of the measurability of tidal effects and the ability of these observatories to constrain the EoS with signals from black hole-neutron star and binary neutron star systems can be found in Refs. Hot+2016; Dam+2012; Del+2013; Aga+2015; Wade+2014 and references therein.
Next, we want to explore the sensitivity of $\Lambda$ to the high-density parametrization of the speed of sound discussed in Sec. III.2. For that purpose, we employ the four EoS used for Fig. 2, keeping the same color and pattern conventions. As for the M(R) results on the left side of Fig. 2, we see that the green curves are what we would expect from the non-conformal calculation and an adiabatic index that’s between the other two, while the speed of sound predictions with or without the conformal limit are dramatically different, see right side of Fig. 2. This seems to indicate that the tidal deformability is mostly sensitive to the EoS at medium to moderately high densities, a regime that “feels” the influence of the microscopic theory S+A_2025, as discussed in Sec. III.1.
@@ IV.2.2 The mass-weighted tidal deformability
with $q=M_{1}/M_{2}$. The chirp mass is one of the most tightly constrained properties extracted from GW170817. Therefore, $M_{c}$ can be set equal to the observed value for the purpose of comparing predictions of $\tilde{\Lambda}$ with experimental data. The GW170817 constraint was stated as $M_{c}/M_{\odot}=1.188^{+0.004}_{-0.002}$ at the 90% confidence level Abb17a, from which the primary and secondary masses (in units of $M_{\odot}$), were inferred to be within the range (1.36, 1.60) and (1.17, 1.36), respectively, placing $q$ between 0.73 and 1. We exploited this information to calculate the values in Table 3. Starting with $M_{1}$ in the full range of masses inferred from GW170817, (1.17, 1.60), and imposing the very stringent $M_{c}$ constraint, we obtain $\tilde{\Lambda}$ for each value of $q$, see Table 3. The effective tidal deformability changes by about 4% across the $q$-variations. Essentially, $\tilde{\Lambda}$ is constrained at fixed mass.
@@
As we already pointed out when discussing $\Lambda_{1.4}$, one of the most physically interesting aspect of the GW170817 data is the opportunity to extract constraints on the radius of the canonical mass neutron star and thus discriminate among various EoS. In Fig. 7, the shaded area represents constraints from GW170817 and AT2017gfo Cough+18, which confine $\tilde{\Lambda}$ to the interval (197, 720). The curve is a typical power law parametrization of $\tilde{\Lambda}$ as a function of $R_{1.4}$, see Ref. Kol+21 and references therein. Clearly, radii smaller than 10.8 km and larger than about 13 km are ruled out.
Figure 7: Effective deformability vs. the radius of the canonical mass neutron star. The shaded area represents constraints from GW170817 and AT2017gfo Cough+18. The continuous curve is a parametrization of $\tilde{\Lambda}$ vs. $R_{1.4}$ Kol+21. Our predictions are indicated by the squares (q=0.73) and the crosses (q=1.0).
@@ V Conclusions and outlook
As noted in Ref. S+A_2025, the maximum-mass constraint moving to higher values, together with the causality requirement at any central density, poses significant restrictions on the high-density continuation of the EoS. Based on those considerations, we selected a small number of EoS differing in the chiral order and/or the adiabatic index of their polytropic portion, followed by a parametrization of the pressure vs. density that garantees causality. We also used a parametriztion where the speed of sound approaches the conformal limit at infinity and observed no significant effect.
We focused on the tidal deformability of neutron stars and related parameters, such as the tidal Love number. Our predictions are consistent with multimessenger constraints, as well as terrestrial laboratory constraints.
We addressed the effective tidal deformability of the binary and the implication of a very tight constraint on the chirp mass of the system. Measurement of the effective tidal deformability (at nearly fixed mass) can be used to extract constraints on the neutron star radius. This is very valuable, as precise neutron star radii would discriminate between equations of state better than any other single measurement. The upper limit on the effective tidal deformability from GW170817 excludes radii larger than 13.2 km, essentially independent of the assumed masses for the component stars. This rules out the findings from PREX II and any steep EoS that generates large radii. We underline the importance of EoS-independent measurements, which can help interpret divergent outcomes from terrestrial experiments that rely on EoS modeling assumptions.

### v2

@@ IV.1 General aspects
The tidal deformability is a parameter that quantifies how easily the star is deformed by an external tidal field. Therefore, a largetidal deformability signals a larger and less compact star which deforms easily.
@@
In the Newtonian limit, the perturbing tidal field, $\epsilon_{ij}$, is the second spacial derivative of the external field, resulting in units of inverse length squared.The quadrupole moment has units of length cubed. We expect $\lambda_{t}$ to be proportional to the fifth power of the neutron star radius, $R$. Adopting traditional convention, we express the tidal deformability as
@@
where C is the previously defined neutron star’s compactness. It is $\Lambda$ that actually governs the effects of tides on inspiral waveforms Van Oeveren and Friedman (2017).
Given a proposed equation of state, the tidal deformability of a neutron star of a certain mass can be obtained by computing the metric in the asymptotic regime using the Einstein equations, extracting the corresponding $r$-order terms, and taking their ratio. The ensuing calculation is described in detail in Ref. Hinderer (2008). The result is an expression for the gravitational Love number, $k_{2}=k_{2}(C,y_{R})$, with $y_{R}$ the solution at the star’s surface of the first order differential equation
@@
and then used in Eq.( 16) to determine the tidal deformability.
Figure 3: Left: The dimensionless tidal deformability as a function of the neutron star mass (left) and as a function of the compactness (right). The pink dots and their uncertainties are taken from the Bayesian analysis of Ref. Lim and Holt (2019). See text for more details.
@@ IV.2.1 Tidal deformability and tidal Love number
Closely related to $\Lambda$ is the tidal Love number $k_{2}$, related to $\Lambda$ as in Eq. (16). This is shown in Fig. 4 as a function of the neutron star mass and compactness. Often, comparison between relativistic and Newtonian calculations or tabulations of deformation properties for polytropic EoS with different polytropic indices are given in terms of $k_{2}$ rather than $\Lambda$ Hinderer (2008). Our predicted values of $k_{2}$ are shown in Table 2 for a representative case.
@@
Recent multimessanger constraints can be found in Ref. Huang (2025), obtained from various methodologies based on variants of the GW170817 data sets. Their most robust constraint is based on a combination of GW-based findings and EoS-independent X-ray constraints, and is reported as
@@
for the dimensionless tidal deformability and radius of the canonical mass neutron star. We underline that this EoS-independent constraint is yet another confirmation that the large radius extracted from the PREX neutron skin measurement Reed et al. (2021) is unrealistic.
We account for both chiral truncation error (cf. Sec. II.1), and the spreading due to different values of $\gamma$. Combining the two independent uncertainties in quadrature, we estimate our ranges for $\Lambda_{1.4}$ and $R_{1.4}$ at N2LO to be
@@
quite consistent with Eq. (21). We underline that the overall uncertainties in Eq. (22) are evaluated at N2LO, which, at this time, is a fully consistent order, see comments in Sec. II.1.
Other inferred upper limits on neutron star radii and tidal deformabilities from GW170817 rule out stiff equations of state that result in large radii and large tidal deformabilities. Recent estimates of the measurability of tidal effects and the ability of these observatories to constrain the EoS with signals from black hole-neutron star and binary neutron star systems can be found in Refs. Hotokezaka et al. (2016); Damour et al. (2012); Del Pozzo et al. (2013); Agathos and et al. (2015); Wade and et al. (2014) and references therein.
Next, we want to explore the sensitivity of $\Lambda$ to the high-density parametrization of the speed of sound discussed in Sec. III.2. For that purpose, we employ the four EoS used for Fig. 2, keeping the same color and pattern conventions. As for the M(R) results on the left side of Fig. 2, we see that the green curves are what we would expect from the non-conformal calculation and an adiabatic index that’s between the other two, while the speed of sound predictions with or without the conformal limit are dramatically different, see right side of Fig. 2. This seems to indicate that the tidal deformability is mostly sensitive to the EoS at medium to moderately high densities, a regime that “feels” the influence of the microscopic theory Sammarruca and Ajagbonna (2025), as discussed in Sec. III.1.
@@
To make the above statement more quantitative, next we perform tests adopting different extension methods. Prior to exploring other options, our typical high-density extension consisted of piecewise polytropes with appropriately chosen adiabatic index. Equations of state unable to support a maximum mass of at least 2.01 M were discarded, and M(R) relations were cut at the central density where causality was violated. With these guidelines, we selected a few favorable combinations, given in Table 3.
Table 3: Maximum mass and $R_{1.4}$ generated by EoS continuations in terms of two polytropes with index $\Gamma_{1}$ and $\Gamma_{2}$, respectively.
EoS label $\Gamma_{1}$ $\Gamma_{2}$ $M_{max}/M_{\odot}$ $R_{1.4}$ (km)
A 3.1 2.7 2.10 12.00
B 3.1 2.8 2.12 12.00
C 3.2 2.7 2.15 12.06
D 3.2 2.8 2.17 12.06
E 3.3 2.7 2.19 12.11
In Table 4, we will compare Love numbers and $\Lambda_{1.4}$ for diverse high-density continuations.
Table 4: Maximum mass and $R_{1.4}$ generated by EoS continuations in terms of two polytropes with index $\Gamma_{1}$ and $\Gamma_{2}$, respectively.
EoS description $k_{2}$ $\Lambda_{1.4}$
A 0.083812 373.908325
B 0.083812 373.908325
C 0.085187 388.95812
D 0.085187 388.95812
E 0.086453 402.891215
N3LO, $\gamma$=3.3 0.086512 403.520110
$(v_{s}/c)^{2}$ $\rightarrow$ 1
N3LO, $\gamma$=3.8 0.091063 457.909912
$(v_{s}/c)^{2}$ $\rightarrow$ 1
N3LO, $\gamma$=3.3 0.086512 403.523460
($v_{s}/c)^{2}$ $\rightarrow$1/3
N3LO450, $\gamma$ = 3.8 0.091064 457.911549
($v_{s}/c)^{2}$ $\rightarrow$ 1/3
For the first group in Table 4, above the first horizontal line, the mild, if any, sensitivity to the high-density continuation is perhaps not surprising, since EoS A through E are not fundamentally different. The second group highlights the sensitivity to the polytropic index employed at medium-high densities. Finally, the third and last group, together with the second, highlights the sensitivity to the super-high density parametrization (for equal $\gamma$). We conclude that, in all of the above cases, differences are essentially negligible compared to the current uncertainties for these quantities. In particular, the last two entries in Table 4 demonstrate that the super-high density continuation has no impact on these observables.
@@ IV.2.2 The mass-weighted tidal deformability
with $q=M_{1}/M_{2}$. The chirp mass is one of the most tightly constrained properties extracted from GW170817. Therefore, $M_{c}$ can be set equal to the observed value for the purpose of comparing predictions of $\tilde{\Lambda}$ with experimental data. The GW170817 constraint was stated as $M_{c}/M_{\odot}=1.188^{+0.004}_{-0.002}$ at the 90% confidence level Abbott and others (2017), from which the primary and secondary masses (in units of $M_{\odot}$), were inferred to be within the range (1.36, 1.60) and (1.17, 1.36), respectively, placing $q$ between 0.73 and 1. We exploited this information to calculate the values in Table 5. Starting with $M_{1}$ in the full range of masses inferred from GW170817, (1.17, 1.60), and imposing the very stringent $M_{c}$ constraint, we obtain $\tilde{\Lambda}$ for each value of $q$, see Table 5. The effective tidal deformability changes by about 4% across the $q$-variations. Essentially, $\tilde{\Lambda}$ is constrained at fixed mass.
@@
As we already pointed out when discussing $\Lambda_{1.4}$, one of the most physically interesting aspect of the GW170817 data is the opportunity to extract constraints on the radius of the canonical mass neutron star and thus discriminate among various EoS. In Fig. 6, the shaded area represents constraints from GW170817 and AT2017gfo Coughlin and others (2018), which confine $\tilde{\Lambda}$ to the interval (197, 720). The curve is a typical power law parametrization of $\tilde{\Lambda}$ as a function of $R_{1.4}$, see Ref. Koliogiannis et al. (2021) and references therein. Clearly, radii smaller than 10.8 km and larger than about 13 km are ruled out.
Figure 6: Effective deformability vs. the radius of the canonical mass neutron star. The shaded area represents constraints from GW170817 and AT2017gfo Coughlin and others (2018). The continuous curve is a parametrization of $\tilde{\Lambda}$ vs. $R_{1.4}$ Koliogiannis et al. (2021). Our predictions are indicated by the squares (q=0.73) and the crosses (q=1.0).
IV.2.3 Additional discussion
Our predictions are generally consistent with multimessenger constraints, as well as terrestrial laboratory constraints. For the purpose of rendering the paper more selfcontained, this short section will address recent terrestrial constraints.
For the symmetry energy, its density dependence, and the pressure in NM, vis a vis constraints from Ref. Lynch and Tsang (2022), the reader is referred to Ref. Sammarruca (2023). Reference Sammarruca (2024) presents a broad analysis of microscopic predictions versus empirical constraints for the neutron skin of 48Ca and 208Pb. In that work, we emphasize that the small skin of 48Ca extracted from CREX is quite consistent with microscopic predictions and available experimental constraints, see Fig.1 of Ref. Sammarruca (2023) and Ref. Sammarruca (2022).
The statistical analysis of Ref. Koehn and et al. (2025) employed 100,000 candidate EoS to obtain Bayesian posteriors on the symmetry energy parameters at normal density. The authors reanalyzed the results of PREX and CREX and concluded that neutron star radius posteriors can be statistically compatible with PREX and CREX. They discuss the question of how each constraint on the EoS is constructed from the observed data, and the various assumptions that could introduce biases. They comment on potential sources of systematic errors while applying a particular constraint, and discuss how different statistical interpretations of the data may affect the inferred results. The authors conclude “… the need for enhanced theoretical models in order to reduce the systematic uncertainties currently present and possibly accommodate both the PREX-II and CREX results.” We take the opportunity to add that theory improvements should never be directed at accomodating specific constraints.
Recently, new constraints Ciampi and et al. (2025) were obtained from the isospin transport ratio in 58,64Ni + 58,64Ni collisions at 32 MeV/nucleon, where the impact parameter dependence was experimentally measured. We understand that the BUU transport model employs effective nuclear energy density functionals that respect the present constraints from ab initio calculations. Figure 7 leaves little doubt on the “outlier” nature of the PREX-II result.
[…]
@@ V Conclusions and outlook
As noted in Ref. Sammarruca and Ajagbonna (2025), the maximum-mass constraint moving to higher values, together with the causality requirement at any central density, poses significant restrictions on the high-density continuation of the EoS. Based on those considerations, we selected a small number of EoS differing in the chiral order and/or the adiabatic index of their polytropic portion, followed by a parametrization of the pressure vs. density that guarantees causality. We also used a parametrization where the speed of sound approaches the conformal limit at infinity and observed no significant impact on the M(R) relation or neutron star deformabilities.
We focused on the tidal deformability of neutron stars and related parameters, such as the tidal Love number. We addressed the effective tidal deformability of the binary and the implication of a very tight constraint on the chirp mass of the system. Measurement of the effective tidal deformability (at nearly fixed mass) can be used to extract constraints on the neutron star radius. This is very valuable, as precise neutron star radii would discriminate between equations of state better than any other single measurement. The upper limit on the effective tidal deformability from GW170817 excludes radii larger than 13.2 km, essentially independent of the assumed masses for the component stars. This rules out the findings from PREX II and any steep EoS that generates large radii. We underline the importance of EoS-independent measurements, which can help interpret divergent outcomes from terrestrial experiments that rely on EoS modeling assumptions. Most recent constraints Tang et al. (2015); Cui and et al. (2025); Ciampi and et al. (2025), from both neutron star observations and heavy ion collisions at Fermi energy, confirm the soft nature of the nuclear equation of state.